# Adaptive Rank, Reduced Forgetting: Knowledge Retention in Continual Learning Vision-Language Models with Dynamic Rank-Selective LoRA

### v6

@@ B.2 Effects of Additional Adaptive Prediction Head on Visual Representations
B.3 Analysis of Changes/Interference w.r.t. Pre-Trained Weights
In Sec. 4.3 of the main paper, Fig. 6 and Fig. 6 visualize and analyze the dynamic ranks assigned to each module. Here, we extend the analysis by comparing parameter changes to pre-trained weights using the amplification factor as in (Hu et al., 2021). The amplification factor is given by:
@@
A higher amplification factor indicates that low-rank parameter updates amplify directions less emphasized in the pre-trained weights (Hu et al., 2021). We visualize the amplification factors for both our method and vanilla LoRA trained on the Aircraft, EuroSAT and Oxford Pets datasets in Fig. 9.
@@
Note that this analysis differs from the visualizations of number of activated ranks in Fig. 6 and Fig. 6. The number of activated ranks reflects the quantity of significant ranks contributing to parameter updates for leaning each task. In this analysis, the amplification factor directly measures the parameter change to each pre-trained weight matrix.
@@
B.4 More Visualizations and Analyses on Rank allocation results of different datasets
@@
B.5 Evaluation with Other VLMs
To evaluate the broader applicability of our method, we conducted preliminary experiments with BLIP (Li et al., 2022) on X-TAIL (Table 9). Our method consistently outperformed prior works and exceeded the pre-trained model’s upper-bound performance, demonstrating its general effectiveness across different VLMs.
@@ B.5 Evaluation with Other VLMs
B.6 Evaluation on Standard Parameter-Efficient Fine-Tuning (PEFT) Task

### v7

@@ B.5 Effects of Additional Adaptive Prediction Head on Visual Representations
B.6 Analysis of Changes/Interference w.r.t. Pre-Trained Weights
In Sec. 4.2 of the main paper, Fig. 6 and Fig. 6 visualize and analyze the dynamic ranks assigned to each module. Here, we extend the analysis by comparing parameter changes to pre-trained weights using the amplification factor as in [27]. The amplification factor is given by:
@@
A higher amplification factor indicates that low-rank parameter updates amplify directions less emphasized in the pre-trained weights [27]. We visualize the amplification factors for both our method and vanilla LoRA trained on the Aircraft, EuroSAT and Oxford Pets datasets in Fig. 9.
@@
Note that this analysis differs from the visualizations of number of activated ranks in Fig. 6 and Fig. 6. The number of activated ranks reflects the quantity of significant ranks contributing to parameter updates for learning each task. In this analysis, the amplification factor directly measures the parameter change to each pre-trained weight matrix.
@@
B.7 Parameter-space Directional Analysis
To further investigate how CoDyRA structures its parameter updates across sequential tasks, we conduct a parameter-space directional analysis following prior work on directional interference in LoRA-based continual learning. Specifically, we compute the cosine similarity between the LoRA $\mathbf{A}$-matrix updates obtained after each task, comparing every task’s update direction against those from all other tasks. This metric captures the degree to which updates for different tasks align or interfere: higher similarity indicates that tasks induce overlapping update directions, while lower similarity reflects more orthogonal and thus less interfering modifications. As shown in Figure 10, CoDyRA consistently produces lower cross-task similarity than a baseline that trains a new LoRA adapter per task and merges them without additional constraints. Despite imposing no explicit orthogonality regularization, CoDyRA naturally encourages more decorrelated update directions, revealing an inherent structural property that helps mitigate destructive interference between tasks.
B.8 More Visualizations and Analyses on Rank allocation results of different datasets
@@
Figure 10: Parameter-space directional analysis of $\mathbf{A}$-matrix updates. For each task, the column displays the average cosine similarity between the model’s LoRA updates after a specific task and the LoRA updates from models trained on all other tasks. This analysis quantifies directional interference among sequential updates. Lower cosine similarity reflects more orthogonal (less interfering) updates. CoDyRA produces lower similarity than the simple LoRA–merging baseline, demonstrating stronger implicit orthogonality of parameter updates even without explicit orthogonal regularization.
@@
B.9 Evaluation with Other VLMs
To evaluate the broader applicability of our method, we conducted preliminary experiments with BLIP [37] on X-TAIL (Table 11). Our method consistently outperformed prior works and exceeded the pre-trained model’s upper-bound performance, demonstrating its general effectiveness across different VLMs.
B.10 Evaluation on Standard Parameter-Efficient Fine-Tuning (PEFT) Task